# Can Large Language Models be Effective Online Opinion Miners?

### v1

@@ Abstract
The surge of user-generated online content presents a wealth of insights into customer preferences and market trends. However, the highly diverse, complex, and context-rich nature of such content poses significant challenges to traditional opinion mining approaches. To address this, we introduce Online Opinion Mining Benchmark (OOMB), a novel dataset and evaluation protocol designed to assess the ability of large language models (LLMs) to mine opinions effectively from diverse and intricate online environments. OOMB provides extensive (entity, feature, opinion) tuple annotations and a comprehensive opinion-centric summary that highlights key opinion topics within each content, thereby enabling the evaluation of both the extractive and abstractive capabilities of models. Through our proposed benchmark, we conduct a comprehensive analysis of which aspects remain challenging and where LLMs exhibit adaptability, to explore whether they can effectively serve as opinion miners in realistic online scenarios. This study lays the foundation for LLM-based opinion mining and discusses directions for future research in this field [^1] https://github.com/ryang1119/Online-Opinion-Mining.
@@
Ryang Heo    Yongsik Seo    Junseong Lee    Dongha Leethanks:   Corresponding author Yonsei University {ryang1119, ysseo, brulee, donalee}@yonsei.ac.kr
@@ 1 Introduction
ASTE Peng et al. (2019) 1,468 15.7 1.7 (a, o, s) Reviews
ACOS Cai et al. (2021) 1,399 15.2 1.5 (a, c, o, s) Reviews
ASQP Zhang et al. (2021) 1,081 14.9 1.5 (a, c, o, s) Reviews
DiaASQ-EN Li et al. (2022) 100 179.7 8.5 (t, a, o, s) Conversation
OOMB (Ours) 600 648.7 14.4 (e, f, o) Reviews, Blogs, Conversation
@@
(2) Confinement to extraction-centric tasks: As mentioned earlier, most prior tasks have focused on extracting opinion spans or structured tuples from input texts. However, this extraction-centric approach can excessively simplify or compress the nuanced contextual information and emotional nuances that are essential for strategic decision-making. For instance, the tuple (“Tesla Model 3”, “interior”, “larger than the previous model”) fails to capture critical contextual background—such as whether the user inspected the vehicle in person or harbored an implicit purchase intent. In real-world industry settings, marketers and product teams are more interested in cohesive, topic-level insights rather than isolated fragments of information Yuan et al. (2015); Santos and Gonçalves (2021); Han et al. (2023b). These observations highlight the need to explore opinion mining paradigms that move beyond raw extraction and aim to preserve the emotions, contextual subtleties, and user intent embedded in real-world discourse.
To address these challenges, we propose Online Opinion Mining Benchmark, named OOMB, a novel benchmark specifically tailored to evaluate the opinion mining capabilities of LLMs across realistic, complex, and diverse online scenarios. Unlike previous datasets, OOMB incorporates content from structurally distinct platforms—including blogs, review sites, Reddit threads, and YouTube comments—capturing long-form content, single & multi-user interactions representative of authentic online discussions. Each content instance is enriched with dual-layer annotations: (1) structured sets of (entity, feature, opinion) tuples reflecting explicit user perspectives, and (2) context-rich, opinion-centric summaries organized around key thematic insights from a marketer’s viewpoint.
@@
  • We present OOMB, a realistic and richly annotated benchmark that evaluates LLMs across structurally diverse online content using both structured tuples and insight-oriented summaries.
@@ 2.2 Data Annotation
For each collected user-generated content, we construct a dual-layer annotation for each content, consisting of both structured opinion tuples and free-form opinion-centric summaries. Following recent studies demonstrating that LLMs with advanced reasoning capabilities can serve as effective tools for data annotation He et al. (2024); Tan et al. (2024), we adopt a human-in-the-loop process in which LLM is first used as the initial annotator, and human annotators verify and refine them to ensure high-quality, reliable labels. The detailed annotation process is described in Appendix A.2.
@@
We annotate each content with structured (entity, feature, opinion) tuples that capture user perspectives. In contrast to the commonly used (aspect, category, opinion, sentiment) schema, this design more closely reflects how real users express opinions—typically without explicit category or sentiment labels, but rather through direct mentions of entity features. Specifically, entity refers to the specific subject or object under discussion in the content (e.g., “Volvo XC90”). feature indicates a characteristic, attribute, or component of the entity that a user mentions or evaluates (e.g., “interior design”). opinion represents the subjective or objective judgment, reaction, experience, evaluation, or feedback regarding a feature (e.g., “luxurious”). If a feature is implicit and does not appear explicitly in the text, it is labeled as “NULL” following Cai et al. (2021). In all other scenarios, each component of the tuple is assumed to be an explicitly mentioned span in the content.
@@
To maximize the coverage and diversity of (entity, feature, opinion) tuple sets from each input content, we perform five rounds of zero-shot prompting using GPT-4o-mini. We then take the union of all generated tuples and remove duplicates to form a distinct preliminary tuple sets. Subsequently, five trained human annotators review every candidate tuple for correctness, eliminate hallucinated entries, and complement any missing tuples. To support consistent decision-making, we design detailed task-specific annotation guidelines and conduct a one-week training session for all annotators, including case-based instruction and edge-case discussions. This process was applied to every content, thereby ensuring high coverage, consistency, and reliability in the final annotations.
Opinion-centric summary
We annotate each content with an opinion-centric summary, a free-form text that organizes diverse opinions into high-level topics for meaningful insights. Specifically, from a marketing manager’s perspective, opinions are grouped into broad categories, highlighting frequently mentioned or standout aspects to reveal key trends. This summary follows a three-to-five-line form, providing a cohesive structure for clear and concise representation of core discussions.
Summary annotation
We generate five independent candidate summaries using the input content and the associated final set of tuples as input. Then, the same five human annotators review each summary from the perspective of a marketing manager and select the highest-quality one that best captures the opinions in the final tuples at the topic level. Similar to the tuple annotation process, we design detailed annotation guidelines to ensure consistent decision-making, and all five annotators undergo a one-week training session. If all candidate summaries are deemed insufficient in quality, the annotators collaboratively rewrite a new summary that more accurately reflects the key insights. For the selected summary, the annotators collaboratively refine and finalize it by checking for missing opinions, eliminating hallucinations, and ensuring conciseness in a three-to-five-line format.
@@ 3 Experiments
3.1 Feature-centric opinion extraction (FOE)
Problem formulation
Evaluation protocol
where $t_{p}=(e_{p},f_{p},o_{p})$ and $t_{g}=(e_{g},f_{g},o_{g})$ are the predicted and gold tuples, respectively. Drawing on recent works Han et al. (2023a); Li et al. (2024), we utilize the Python’s difflib library [^3] https://docs.python.org/3/library/difflib.html to compute token-level overlap scores for lexical similarity (L-RM), while employing a Sentence Transformer [^4] We use all-MiniLM-L6-v2, a lightweight model optimized for efficient sentence similarity computation. for semantic similarity (S-RM). (3) Contextual Match (CM): Inspired by Fu et al. (2023); Fane et al. (2025), we design a method that leverages the reasoning capabilities of LLMs to match tuples in a manner similar to human judgment. Specifically, we utilize GPT-4o to evaluate both predicted and gold tuples, enabling the model to count how many tuples match. This metric allows recognition of semantically equivalent tuples even when surface forms differ significantly, used the prompt shown in Table 17. Note that for both RM and CM, we measure recall by counting each gold tuple at most once to avoid double counting, even if multiple predicted tuples match the same gold tuple. For all evaluation metrics, we primarily use the F1 score while also reporting precision and recall.
@@ 3.2 Opinion-centric insight generation (OIG)
This task aims to analyze whether LLMs can group scattered opinions from user-generated online content into high-level topics, providing context-aware and meaningful insights. Formally, given content $c$, our objective is to generate a free-form text summary $S$ that cohesively encapsulates user opinions into high-level topics.
Evaluation protocol
To broadly assess the quality of opinion insight summaries generated by the model across various aspects, we employ both lexical and semantic automated evaluation metrics. For lexical evaluation, we adopt ROUGE-1,10, L Lin (2004), which measure word overlap between the reference and generated summaries. For semantic evaluation, we leverage BERTScore (BS) Zhang et al. (2019) and A3CU Liu et al. (2023b). BS computes similarity between the reference and generated texts using contextual embeddings, while A3CU compares texts without extracting atomic content units, providing a human-aligned assessment of content similarity. For both ROUGE and A3CU, we report F1 scores. Moreover, to ensure a systematic and comprehensive evaluation, we also conduct reference-free assessments using an LLM as the judge. Inspired by Siledar et al. (2024), we design the following six well-defined criteria: Faithfulness, Coverage, Specificity, Insightfulness, Intent and Fluency. This analysis extends beyond automated lexical and semantic metrics, providing a broader perspective on the abstractive opinion mining capabilities of LLMs. A detailed description is provided in Appendix B.3.
3.3 Experimental setup
Models
Implementation details
@@ 4.1 RQ1: What makes it challenging for LLMs to extract structured opinions?
To investigate how extensively an LLM can extract structured tuples from content, we perform multiple inference iterations per single input and measure the model’s extraction performance. For evaluation, we take the union of all tuples generated across iterations, remove duplicates, and consider only the unique (entity, feature, opinion) tuple sets. To capture a broader range of tuple sets, we set the temperature to 1.0 during inference. As shown in Figure 3, most models generate a significantly larger number of predicted tuples as the number of inference iterations increases, but the number of correctly matched tuples does not keep pace. Notably, recall improves significantly across most models but eventually reaches a plateau, where the rate of increase diminishes. This implies that LLMs recognize a fixed set of opinions within the content, making it challenging to cover every opinion merely by increasing the number of inference iterations. Therefore, improving the extraction capabilities of LLMs requires exploring alternative strategies beyond merely repeating the inference process.
@@
4.2 RQ2: How insightfully can LLMs generate abstractive opinion summaries?
Automated evaluation results
Table 5 reports the performance for the OIG task, using both lexical and semantic evaluation metrics. While models show strong word-level overlap (R-1 and R-L) in their summaries, they exhibit significantly lower bigram recall (R-2), highlighting difficulty in sustaining coherent phrase structures. Additionally, they achieve relatively high BS; their performance on A3CU remains substantially lower, suggesting that LLMs often capture surface-level semantic similarity but struggle to reflect deeper, human-aligned content understanding. Thus, to thoroughly gauge LLMs’ abstractive strengths—particularly their capture of intent, subtle sentiment shifts, and deeper insights beyond surface semantics—a multifaceted evaluation framework is needed.
@@ 4.2 RQ2: How insightfully can LLMs generate abstractive opinion summaries?
LLM-Judge evaluation across multiple perspectives
To comprehensively analyze how well models generate abstractive opinion summaries, we conduct a reference-free evaluation using an LLM as the judgemeter. From the results in Figure 4, we derive the following key conclusions: (1) LLMs consistently provide natural and readable summaries while preserving the original content without distortion or unnecessary modification. This demonstrates their strength in faithfulness and fluency, ensuring that the generated summaries remain accurate and coherent. (2) However, LLMs struggle to capture implicit user intentions, nuanced expressions, and meaningful insights that are not explicitly stated in the input content. This limitation is reflected in lower scores for insightfulness and intent, indicating that while LLMs can summarize well, they lack deeper abstraction and contextual understanding.
Impact of structured opinions on summary
Figure 6 demonstrates that augmenting opinion tuples during opinion-centric summary generation not only leads to substantial improvements in automatic evaluation metrics, but also provides practical benefits from a user perspective. In particular, the notable gains in coverage and insightfulness suggest that the model becomes more effective at capturing key opinions and delivering more informative summaries. Conversely, slight decreases in intent and fluency indicate that the added structure may sometimes interfere with natural expression and tone preservation. These results suggest that integrating structured opinion tuples into the insight generation pipeline is a key strategy for effective opinion mining, while also highlighting the need for continued research into both the extractive and abstractive capabilities of LLMs.
Figure 4: Radar charts for LLM-Judge evaluations of the OIG task. Comparison of the average model performance across different content types (Left). Comparison of performance across different models (Right).
@@ 5 Related Work
Recently, large language models (LLMs) OpenAI et al. (2024); Grattafiori et al. (2024); Team et al. (2024a) have demonstrated remarkable zero-shot and in-context learning capabilities across a range of tasks, including information extraction Kim et al. (2024b); Perot et al. (2024); Liu et al. (2024) and abstractive summarization Chhabra et al. (2024); Tang et al. (2024); Siledar et al. (2024). While these advances suggest that LLMs have great potential in opinion mining, existing benchmarks fall short of capturing the complexity of real-world inputs and remain focused on simplified, structured extraction settings. As a result, they fail to fully assess the true potential of LLMs in this domain. To bridge this gap, we introduce the OOMB benchmark, which encompasses a wide spectrum of realistic online content and enables comprehensive investigation of both extraction and abstractive capacities of LLMs.

### v2

@@ Abstract
The surge of user-generated online content presents a wealth of insights into customer preferences and market trends. However, the highly diverse, complex, and context-rich nature of such content poses significant challenges to traditional opinion mining approaches. To address this, we introduce Online Opinion Mining Benchmark (OOMB), a novel dataset and evaluation protocol designed to assess the ability of large language models (LLMs) to mine opinions effectively from diverse and intricate online environments. OOMB provides, for each content instance, an extensive set of (entity, feature, opinion) tuples and a corresponding opinion-centric insight that highlights key opinion topics, thereby enabling the evaluation of both the extractive and abstractive capabilities of models. Through our proposed benchmark, we conduct a comprehensive analysis of which aspects remain challenging and where LLMs exhibit adaptability, to explore whether they can effectively serve as opinion miners in realistic online scenarios. This study lays the foundation for LLM-based opinion mining and discusses directions for future research in this field. Our code and dataset are available [^1] https://github.com/ryang1119/OOMB.
@@
Ryang Heo    Yongsik Seo    Junseong Lee    Dongha Leethanks:   Corresponding author Yonsei University - DLI Lab {ryang1119, ysseo, brulee, donalee}@yonsei.ac.kr
@@ 1 Introduction
ASTE Peng et al. (2019)
1,468 15.7 1.7 (a, o, s) Reviews
ACOS Cai et al. (2021)
1,399 15.2 1.5 (a, c, o, s) Reviews
ASQP Zhang et al. (2021)
1,081 14.9 1.5 (a, c, o, s) Reviews
DiaASQ-EN Li et al. (2023)
100 179.7 8.5 (t, a, o, s) Conversation
OOMB (Ours)
600 648.7 14.4 (e, f, o) Reviews, Blogs, Conversation
@@
(2) Confinement to extraction-centric tasks: As mentioned earlier, most prior tasks have focused on extracting opinion spans or structured tuples from input texts. However, this extraction-centric approach can excessively simplify or compress the nuanced contextual information and emotional nuances that are essential for strategic decision-making. For instance, the tuple (“Tesla Model 3”, “interior”, “larger than the previous model”) fails to capture critical contextual background, such as whether the user inspected the vehicle in person or harbored an implicit purchase intent. In real-world industry settings, marketers and product teams are more interested in cohesive, topic-level insights rather than isolated fragments of information Yuan et al. (2015); Santos and Gonçalves (2021); Han et al. (2023b). These observations highlight the need to explore opinion mining paradigms that move beyond raw extraction and aim to preserve the emotions, contextual subtleties, and user intent embedded in real-world discourse.
To address these challenges, we propose Online Opinion Mining Benchmark, named OOMB, a novel benchmark specifically tailored to evaluate the opinion mining capabilities of LLMs across realistic, complex, and diverse online scenarios. Unlike previous datasets, OOMB incorporates content from structurally distinct platforms, including blogs, review sites, Reddit threads, and YouTube comments, capturing long-form content as well as single & multi-user interactions representative of authentic online discussions. Each content instance is enriched with dual-layer annotations: (1) a structured set of (entity, feature, opinion) tuples reflecting explicit user perspectives, and (2) a context-rich, opinion-centric insight organized around key thematic topics from a marketer’s viewpoint.
@@
  • We present OOMB, a realistic and richly annotated benchmark that evaluates LLMs across structurally diverse online content using both structured tuples and opinion-centric insights.
@@ 2.2 Data Annotation
For each collected user-generated content, we construct a dual-layer annotation for each content, consisting of both structured opinion tuples and free-form opinion-centric insights. Following recent studies demonstrating that LLMs with advanced reasoning capabilities can serve as effective tools for data annotation He et al. (2024); Tan et al. (2024), we adopt a human-in-the-loop process in which LLM is first used as the initial annotator, and human annotators verify and refine them to ensure high-quality, reliable labels. The detailed annotation process is described in Appendix A.2.
@@
We annotate each content with structured (entity, feature, opinion) tuples that capture user perspectives. In contrast to the commonly used (aspect, category, opinion, sentiment) schema, this design more closely reflects how real users express opinions. Typically, they do so without explicit category or sentiment labels but rather through direct mentions of entity features. Specifically, entity refers to the specific subject or object under discussion in the content (e.g., “Volvo XC90”). feature indicates a characteristic, attribute, or component of the entity that a user mentions or evaluates (e.g., “interior design”). opinion represents the subjective or objective judgment, reaction, experience, evaluation, or feedback regarding a feature (e.g., “luxurious”). If a feature is implicit and does not appear explicitly in the text, it is labeled as “NULL” following Cai et al. (2021). In all other scenarios, each component of the tuple is assumed to be an explicitly mentioned span in the content.
@@
To maximize the coverage and diversity of (entity, feature, opinion) tuples from each input content, we perform five rounds of zero-shot prompting using GPT-4o-mini 333gpt-4o-mini-2024-07-18. We then take the union of all generated tuples and remove duplicates to form a distinct preliminary set of tuples. Subsequently, five trained human annotators review every candidate tuple for correctness, eliminate hallucinated entries, and complement any missing tuples. To support consistent decision-making, we design detailed task-specific annotation guidelines and conduct a one-week training session for all annotators, including case-based instruction and edge-case discussions. This process was applied to every content, thereby ensuring high coverage, consistency, and reliability in the final annotations.
Opinion-centric insight
We annotate each content with an opinion-centric insight, a free-form text that organizes diverse opinions into high-level topics for meaningful insights. Specifically, from a marketing manager’s perspective, opinions are grouped into broad categories, highlighting frequently mentioned or standout aspects to reveal key trends. This insight follows a three-to-five-line form, providing a cohesive structure for clear and concise representation of core discussions.
Insight annotation
We generate five independent candidate insights using the input content and the associated final set of tuples as input. Then, the same five human annotators review each insight from the perspective of a marketing manager and select the highest-quality one that best captures the opinions in the final tuples at the topic level. Similar to the tuple annotation process, we design detailed annotation guidelines to ensure consistent decision-making, and all five annotators undergo a one-week training session. If all candidate insights are deemed insufficient in quality, the annotators collaboratively rewrite a new insight that more accurately reflects the key insights. For the selected insight, the annotators collaboratively refine and finalize it by checking for missing opinions, eliminating hallucinations, and ensuring conciseness in a three-to-five-line format.
Task Metric Score
Span-level Tuple Agreement Fleiss’ Kappa 0.6821
Preferred Insight Agreement Fleiss’ Kappa 0.7695
Insight Semantic Consistency BERTScore 0.9297
Table 2: Inter-annotator agreement score across three different annotation stages (p-value < 0.05). Detailed description of the analysis is provided in the Appendix A.2
2.3 Inter-Annotator Agreement Analysis
To further ensure the reliability of our annotation process, we conduct an inter-annotator agreement study. Five trained annotators participated in this analysis across 100 randomly sampled contents. We assess agreement at three different stages of the annotation process: (1) Span-level tuple agreement, (2) Preferred insight agreement, and (3) Insight semantic consistency. Table 2 demonstrate that annotators reached substantial agreement in both structured tuple extraction and insight selection, while finalized insights exhibited strong semantic consistency across annotators. This comprehensive analysis confirms that our benchmark was constructed through a highly reliable process.
@@ 3 Experiments
3.1 Feature-centric opinion extraction (FOE)
Problem formulation
Evaluation metric
where $t_{p}=(e_{p},f_{p},o_{p})$ and $t_{g}=(e_{g},f_{g},o_{g})$ are the predicted and gold tuples, respectively. Drawing on recent works Han et al. (2023a); Li et al. (2024), we utilize the Python’s difflib library [^4] https://docs.python.org/3/library/difflib.html to compute token-level overlap scores for lexical similarity (L-RM), while employing a Sentence Transformer555sentence-transformers/all-MiniLM-L6-v2 for semantic similarity (S-RM). (3) Contextual Match (CM): Inspired by Fu et al. (2024); Fane et al. (2025), we design a method that leverages the reasoning capabilities of LLMs to match tuples in a manner similar to human judgment. Specifically, we utilize GPT-4o 666gpt-4o-2024-08-06 to evaluate both predicted and gold tuples, enabling the model to count how many tuples match. This metric allows recognition of semantically equivalent tuples even when surface forms differ significantly, used the prompt shown in Table 18. Note that for both RM and CM, we measure recall by counting each gold tuple at most once to avoid double counting, even if multiple predicted tuples match the same gold tuple. For all evaluation metrics, we primarily use the F1 score while also reporting precision and recall.
@@ 3.2 Opinion-centric insight generation (OIG)
This task aims to analyze whether LLMs can group scattered opinions from user-generated online content into high-level topics, providing context-aware and meaningful insights. Formally, given content $c$, our objective is to generate a insightful text $I$ that cohesively encapsulates user opinions into high-level topics.
Evaluation metric
To broadly assess the quality of opinion-centric insights generated by the model across various aspects, we employ both lexical and semantic automated evaluation metrics. For lexical evaluation, we adopt ROUGE-1, 10, L Lin (2004), which measure word overlap between the reference and generated insights. For semantic evaluation, we leverage BERTScore (BS) Zhang et al. and A3CU Liu et al. (2023b). BS computes similarity between the reference and generated texts using contextual embeddings, while A3CU compares texts without extracting atomic content units, providing a human-aligned assessment of content similarity. For both ROUGE and A3CU, we report F1 scores. Moreover, to ensure a systematic and comprehensive evaluation, we also conduct reference-free assessments using an LLM-as-a judge. Inspired by Siledar et al. (2024), we design the following six well-defined criteria: Faithfulness, Coverage, Specificity, Insightfulness, Intent and Fluency. This analysis extends beyond automated lexical and semantic metrics, providing a broader perspective on the abstractive opinion mining capabilities of LLMs. A detailed description is provided in Appendix B.3.
@@ 4.1 RQ1: What makes it challenging for LLMs to extract structured opinions?
To investigate how extensively an LLM can extract structured tuples from content, we perform multiple inference iterations per single input and measure the model’s extraction performance. For evaluation, we take the union of all tuples generated across iterations, remove duplicates, and consider only the unique set of (entity, feature, opinion) tuples. To capture a broader range of tuples, we set the temperature to 1.0 during inference. As shown in Figure 3, most models generate a significantly larger number of predicted tuples as the number of inference iterations increases, but the number of correctly matched tuples does not keep pace. Notably, recall improves significantly across most models but eventually reaches a plateau, where the rate of increase diminishes. This implies that LLMs recognize a fixed set of opinions within the content, making it challenging to cover every opinion merely by increasing the number of inference iterations. Therefore, improving the extraction capabilities of LLMs requires exploring alternative strategies beyond merely repeating the inference process.
@@
4.2 RQ2: How meaningfully can LLMs generate meaningful opinion insights?
Automated evaluation results
Table 6 reports the performance for the OIG task, using both lexical and semantic evaluation metrics. While models show strong word-level overlap (R-1 and R-L) in their insights, they exhibit significantly lower bigram recall (R-2), highlighting difficulty in sustaining coherent phrase structures. Additionally, they achieve relatively high BS; their performance on A3CU remains substantially lower, suggesting that LLMs often capture surface-level semantic similarity but struggle to reflect deeper, human-aligned content understanding. Thus, to thoroughly gauge LLMs’ abstractive strengths, particularly in capturing intent, subtle sentiment shifts, and deeper insights beyond surface semantics, a multifaceted evaluation framework is needed.
Figure 4: Radar charts for LLM-as-a-judge evaluations of the OIG task. Comparison of the average model performance across different content types (Left). Comparison of performance across different models (Right).
LLM-as-a-judge evaluation across multiple perspectives
To comprehensively analyze how well models generate abstractive opinion-centric insights, we conduct a reference-free evaluation using an LLM as the judgemeter. From the results in Figure 4, we derive the following key conclusions: (1) LLMs consistently provide natural and readable insights while preserving the original content without distortion or unnecessary modification. This demonstrates their strength in faithfulness and fluency, ensuring that the generated insights remain accurate and coherent. (2) However, LLMs struggle to capture implicit user intentions, nuanced expressions, and meaningful insights that are not explicitly stated in the input content. This limitation is reflected in lower scores for insightfulness and intent, indicating that while LLMs can summarize well, they lack deeper abstraction and contextual understanding.
Impact of structured opinions on insight generation
Figure 6 demonstrates that augmenting opinion tuples during opinion-centric insight generation not only leads to substantial improvements in automatic evaluation metrics, but also provides practical benefits from a user perspective. In particular, the notable gains in coverage and insightfulness suggest that the model becomes more effective at capturing key opinions and delivering more informative opinion-centric insights. Conversely, slight decreases in intent and fluency indicate that the added structure may sometimes interfere with natural expression and tone preservation. These results suggest that integrating structured opinion tuples into the insight generation pipeline is a key strategy for effective opinion mining, while also highlighting the need for continued research into both the extractive and abstractive capabilities of LLMs.
@@ 5 Related Work
Recently, large language models (LLMs) OpenAI et al. (2024); Grattafiori et al. (2024); Team et al. (2024a) have demonstrated remarkable zero-shot and in-context learning capabilities across a range of tasks, including information extraction Kim et al. (2024b); Perot et al. (2024); Liu et al. (2024) and abstractive summarization Chhabra et al. (2024); Tang et al. (2024); Siledar et al. (2024). While these advances suggest that LLMs have great potential in opinion mining, existing benchmarks fall short of capturing the complexity of real-world inputs and remain focused on simplified, structured extraction settings. As a result, they fail to fully assess the true potential of LLMs in this domain. To bridge this gap, we propose the OOMB benchmark, which encompasses a wide spectrum of realistic online content and enables comprehensive investigation of both extraction and abstractive capacities of LLMs.